OpenStudy (zyberg):

How to solve |x^2 - 5x + 4| = 4? (Why is my solution wrong?)

OpenStudy (zyberg):

My solution, that gives off too many roots: First I split this into two cases: 1) x^2 - 5x + 4 = 4 2) -(x^2 - 5x + 4) = 4; x^2 - 5x + 4 = -4 Solving 1: x(x - 5) = 0; x = 0 or x = 5. Solving 2: x^2 - 5x + 8 = 0; x = 2 or 3. The answers in my book say that it's only 0, 5 (so, only the first case is the right one). Why is that?

OpenStudy (solomonzelman):

\(x^2 - 5x + (2.5/2)^2 = -8+ (2.5/2)^2\)

OpenStudy (solomonzelman):

\(x^2 - 5x + (2.5/2)^2 = -8+ (2.5/2)^2\) \(x^2 - 5x + (6.25/4) = -8+ (6.25/4)\) \((x-2.5)^2 =(-32+6.25)/4<0\) See what I am doing?

OpenStudy (solomonzelman):

so, your solutions are clearly imaginary numbers. (No real solutions.)

OpenStudy (zyberg):

Thank you very much! :)

OpenStudy (solomonzelman):

No problem:)